OpenStudy (anonymous):

which expression is an equivalent form of 2x^3 + 4x^2/ 2x^2 show work please F: x+2 G: 2x(x+1) H: 2x^2 (x+1) J: 2x^2(x+3)

OpenStudy (agreene):

F: x+2 \[\frac{2x^3+4x^2}{2x^2}=\frac{2x^3}{2x^2}+\frac{4x^2}{2x^2}=x+2\]